OpenStudy (fools101):

help please!!! Geometry! How do i do this?! Equations of Circles Find the equation of the circle with the given center and radius. C(0, 0), r = 4 A) (x + 4)2 + (y + 4)2 = 4 B) x2 + y2 = 4 C) x2 + y2 = 16 D) (x - 2)2 + (y - 2)2 = 16

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C

OpenStudy (fools101):

thank u but how did u get that answer? Find the equation of the circle with the given center and radius. C(2, -1); r = 3 A) (x + 2)2 + (y - 1)2 = 3 B) (x - 2)2 + (y + 1)2 = 9 C) (x + 2)2 + (y - 1)2 = 9 D) (x - 2)2 + (y + 1)2 = 3

OpenStudy (anonymous):

The equation of a circle is (x-xc)^2 + (y-yc)^2 = r^2 where (xc,yc) are the coordinates of the center and r is the radius. Replace to verify that it is C

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] is general equation for circle with center (h,k) radius r

OpenStudy (anonymous):

in your case you have \[h=2,k=-1,r=9\] so you get \[(x-2)^2+(y+1)^2=9\]

OpenStudy (anonymous):

sorry i meant \[h=2,k=-1,r=3\]

OpenStudy (fools101):

oh so 2 is A?!

OpenStudy (fools101):

I get a little thank u...